If 0.590% solution of Na2CO3 will be used to precipitate Ca+2 from solution, determine the amount of solution, in grams needed to percipitate 5.00x102mL of 0.01120M Ca+2

1
Expert's answer
2022-03-16T10:04:03-0400

V(Ca2+) = 500.0 mL = 0.5 L;

C(Ca2+) = 0.01120 M;

n(Ca2+) = C(Ca2+) * V(Ca2+) = 0.01120 * 0.5 = 0.0056 moles;

Na2CO3 + Ca2+ = CaCO3 + 2Na+;

For the chemical reaction:

n(Na2CO3) = n(Ca2+) = 0.0056 moles;

X(Na2CO3) = 0.590%

M(Na2CO3) = 106 g/mole;

m(Na2CO3) = n(Na2CO3) * M(Na2CO3) = 0.0056 * 106 = 0.5936 g;

m(solution) = m(Na2CO3) * 100%/X(Na2CO3) = 0.5936 * 100/0.590 = 100.6 g.

Answer: 100.6 g.